OpenStudy (anonymous):

What is the value of the discriminant of the polynomial 3x^2+6x+5?

OpenStudy (anonymous):

The discriminant is \[b^2-4ac\]a=3, b=6, and c=5

myininaya (myininaya):

\[b^2-4ac=(6)^2-4(3)(5)=36-60=-24<0, no real solution\]